OpenStudy (anonymous):

2.5 + 2x = 0.5 solve for x

OpenStudy (anonymous):

we always try to get x by itself. so first get rid of the 2.5. do you know how to do thhis part?

OpenStudy (anonymous):

...do I subtract 0.5(or 1/2) from it?

OpenStudy (lstrasmann):

2.5 + 2x = 0.5 -2.5 -2.5 ------------ 2x = -2 2x / 2 = -2 / 2 x = -1

OpenStudy (anonymous):

I checked that. It turns out like this: 2.5 + 2(-1) = 0.5 4.5 = 0.5 it's wrong.

OpenStudy (lstrasmann):

How did you get 4.5? 2.5 + 2(-1) = 2.5 + -2 = 0.5

OpenStudy (anonymous):

I multiplied 2 by -1, then I added it to 2.5 like the eqaution says...

OpenStudy (lstrasmann):

2 x -1 = -2, so 2.5 + -2 = 0.5

OpenStudy (anonymous):

oh! I was doing it wrong. xD haha stupid me. thanks you!

OpenStudy (lstrasmann):

you're welcome! sometimes i do that too ;)

OpenStudy (anonymous):

haha I am not exactly a math wizz